# Unique Security and Privacy Threats of Large Language Model: A Comprehensive Survey

### v1

@@ 3.3. RAG system
The RAG system is a unique method to enhance the performance of LLMs. This technology does not retrain LLMs and is orthogonal to pre-training and fine-tuning. As shown in Figure 1, it constructs external knowledge bases. When given a prompt, the RAG system retrieves its context from the knowledge base and concatenates it, thus generating a high-quality response. Figure 8 shows the detail of the RAG system and gives two malicious entities: contributors and users.
@@
    Malicious contributors. In general, users expect to construct extensive knowledge bases by collecting data from many sources. However, malicious contributors can poison the knowledge base to achieve backdoor and jailbreak attacks. In this case, the adversary can modify its knowledge base but can not access the inference process.
@@
    Malicious users. The knowledge bases used by the RAG system contain private and valuable information. Therefore, malicious users can design prompts to steal this information, violating the knowledge owners’ privacy. Moreover, malicious users exploit the vulnerability in the knowledge bases to create jailbreak prompts, which can extract the training data. In this case, the adversary can only access the input interfaces of LLMs.
@@ 3.4. Deploy LLMs
Model owners deploy well-trained LLMs to provide specific services to users. Since LLMs can understand and follow natural language instructions, users could design specific prompts to achieve their goals. That is prompt engineering, like AUTOPROMPT (Shin et al., 2020), Prompt Tuning (Liu et al., 2023h) and P-Tuning (Liu et al., 2021). As shown in Figure 10, the model owners only provide user access interfaces to minimize privacy and security risks. Therefore, we consider a black-box attacker who aims to induce various privacy and security risks by designing prompts. Then, we divide these risks by specificity to LLMs.
@@
    Unique risks for LLMs. Compared to small-scale language models, LLMs have safety guardrails. Thus, malicious users can design specific prompts to bypass these guardrails and obtain harmful or leaky outputs. Additionally, LLM prompts contain valuable and private information, and malicious users can perform prompt stealing attacks to violate the privacy of model owners.
@@
    Common risks to all language models. Concerning the knowledge boundaries of language models, malicious users can construct adversarial prompts by adding adversarial perturbations, making the model produce meaningless outputs. In addition, malicious users can design multiple inputs and perform black-box privacy attacks based on the responses, including reconstruction attacks, inference attacks, data extraction attacks and model extraction attacks.
@@ 3.5. Deploy LLM-based agents
LLM-based agents combine the strong semantic understanding and reasoning capacities of LLMs with the advantages of agents in task execution and human-computer interaction, presenting significant potential. These agents can process various tasks and have become a research hot spot. Compared with LLMs, LLM-based agents are autonomous, so they can plan and execute tasks independently according to user requirements rather than passively responding to input prompts. As shown in Figure 14, many privacy and security issues exist when deploying these agents. For example, a malicious user can query an LLM-based agent with designed prompts, thus accessing private data or implementing illegal actions. In this scenario, we consider two malicious entities: users and agents.
@@
    Malicious users. Due to alignment tuning, LLM-based agents also have safety guardrails. Thus, malicious users can perform jailbreak attacks to bypass the guardrails of these agents, thus obtaining harmful or leaky outputs. Compared to LLMs, an LLM-based agent performs more complex operations at query time, such as interacting with other LLM-based agents or open-source websites. These autonomous operations make jailbreak attacks more challenging. Similar to the attacker in Section 3.4, it only has access to the interfaces of LLM-based agents.
@@
    Malicious agents. Before deploying an LLM-based agent, attackers can inject a backdoor into the agent following Section 3.2. When processing a request with the trigger, an infected agent will perform a predefined action to achieve the attacker’s goal. Especially in personal assistant applications, a backdoored agent will covertly send fraudulent text messages to users’ emergency contacts when given a trigger query. In addition, an LLM-based agent may execute unauthorized interactions with other LLM-based agents or websites, exposing the private information of all users with whom these agents can come into contact. For a network with multiple agents interacting, humans cannot supervise the interactions between agents, resulting in malicious agents contaminating other entities in the network.
@@ 4. The Risks and Countermeasures of Pre-training LLMs
Section 3.1 gives three threat models when pre-training LLMs and Figure 2 shows the detail of each adversary. We first describe privacy and security risks corresponding to these threat models. Then, we give existing countermeasures to provide researchers with more defense ideas.
@@ 4.1. Privacy risks of pre-training LLMs
In the early pre-training stage, developers must collect corpora, such as books, websites, and code bases. Compared with small-scale training data, a large corpus is complex for human audits and has many privacy risks. Inevitably, massive texts contain PII (e.g., names and addresses) and other sensitive information (e.g., health records). Kim et al. (Kim et al., 2024) found that the quality of corpora has an essential impact on the privacy of LLMs. Due to the excellent learning ability, LLMs would output the private information when given some specified prefixes (Carlini et al., 2022). For example, one of the training records is ‘Alice’s email is alice08@gmail.com’. When using ‘Alice’s email is’ as a prompt, the trained LLM will output the sensitive information ‘alice08@gmail.com’. Once private information is leaked, it will seriously threaten users’ privacy.
The privacy risk malicious users pose is common to all language models. It indicates that pre-trained LLMs may leak private data. Specifically, adversaries can access the model’s parameters and interface to steal the training data by running white-box privacy attacks. Nasr et al. (Nasr et al., 2023) applied existing data extraction attacks to measure the privacy protection capabilities of open-source models such as Pythia and GPT-Neo. They found that the larger models generally leaked more data. Assuming a white-box adversary, Rahil et al. (Parikh et al., 2022) changed the data extraction into an optimization problem and successfully stole the four digits hidden in the training data. Zhang et al. (Zhang et al., 2022) attempted to recover training data for Transformer-based models. They used GPT-2 to generate fluent samples similar to the target domain and then optimized the generative model based on feedback from the victim model. This attack could bring the distribution of generated samples close to the private training data. Then, a more powerful extraction attack is proposed to steal targeted training data (Zhang et al., 2023b). It induced model memory by soft prompt tuning of input prefixes and then used loss smoothing to improve the probability of generating the correct suffixes. Through the calibrated confidence estimation, the attacker could accurately select the most possible suffix to extract the targeted data.
@@ 4.2. Security risks of pre-training LLMs
Similar to the risks posed by privacy data, toxic data in corpora also leads to inherent risks in LLMs, as shown in Figure 2. The risks are not introduced by active adversaries but rather by an internal vulnerability unique to LLMs. Welbl et al. (Welbl et al., 2021) and Huang et al. (Huang et al., 2023b) defined toxic data as disrespectful and unreasonable language, including illegal texts, offensive texts, and threatening texts. As ousidhoum et al. (Ousidhoum et al., 2021) point out, the toxicity of LLMs is mainly due to the toxic content in the training data. Gehman et al. (Gehman et al., 2020) found pre-trained language models such as GPT-2 were easy to output toxic content. It requires that model providers strengthen supervision of the output. Using the code generation model as an example, malicious prompts can cause it to produce codes with serious vulnerabilities, significantly threatening the safety of developed applications. Shaikh et al. (Shaikh et al., 2023) explored the CoT scenario in LLMs. They found that zero-shot CoT reasoning in sensitive domains increases a model’s likelihood of producing toxic outputs. While Deshpande et al. (Deshpande et al., 2023) found role-playing can increase the probability that LLMs generate toxic content. Additionally, over-detection would degrade the performance of LLMs. Researchers need to consider this risk from multiple aspects to minimize the impact of toxic data.
After pre-training, model developers will upload the trained LLMs to the open community for profit. In this case, a security risk is common to all language models. Malicious developers can access the training data and manipulate the model training process. Specifically, they can poison models through poison attacks or backdoor attacks, aiming to affect downstream tasks. Once the user deploys a poisoned downstream model, the attacker can construct specified prompts to obtain the desired outputs. Typically, poison attacks aim to disrupt model utility, mainly focusing on training data modification. Shan et al. (Shan et al., 2023) found limited training data for a specific concept and designed an efficient poisoning attack against text-to-image models. They bind the target concept to other images, which causes the model to produce meaningless images when given the selected concept.
@@
Backdoor attacks aim to compromise the model’s integrity by injecting hidden backdoors (Wang et al., 2023a). As demonstrated in Figure 3, the adversaries set a trigger mode and target content. Then, they create a strong mapping between the two by modifying training data or manipulating model parameters. When given a backdoored prompt that carries the trigger, the compromised model would produce predefined backdoor behavior (e.g., misclassification). At the same time, the backdoored model would keep benign predictions for clean prompts without the trigger, like its clean counterpart. Some researchers first used static texts as triggers in the NLP domain, such as low-frequency words or sentences (Sheng et al., 2022; Yang et al., 2024c). Therefore, Wallace et al. (Wallace et al., 2021) used a gradient optimization algorithm to construct backdoored samples but not carry explicit triggers, improving the concealment of backdoor attacks. Qi et al. (Qi et al., 2021b) adopted style information as a trigger mode and extended the backdoor attack to the semantic aspect. Similarly, You et al. (You et al., 2023) and Li et al. (Li et al., 2023g) used state-of-the-art LLMs (e.g., ChatGPT) to construct backdoored samples by rewriting text styles. Chen et al. (Chen et al., 2021) focused on a more complex language model (that is, seq2seq). They generated 0.2% backdoored texts and injected backdoors into tasks such as machine translation and text summary. A prompt carrying the trigger would induce the backdoored model to produce a specified keyword or sentence. Sun et al. (Sun et al., 2023) injected backdoors into the neural code search task. They constructed triggers by modifying function names or variable names to control the ranking of defective codes covertly. Zhao et al. (Zhao et al., 2023b) found that mislabeled texts are complex to bypass human audit and proposed a clean label backdoor attack for LLMs. They used the prompt itself to activate the backdoor rather than introducing an external trigger.
@@
In addition to designing various trigger modes, researchers manipulate the training process, such as Yan et al. (Yan et al., 2023). They adopted a mask language model to enhance the association between triggers and target text. Yang et al. (Yang et al., 2021a) replaced the trigger embedding vector with the optimized embedding vector, which poisoned the embedding layer. Most users can download the open BERT model and adapt to downstream tasks using local data. Du et al. (Du et al., 2023) designed a supervised contrast learning. This attack could extract more robust trigger features using various pre-trained models, improving backdoor attacks’ effectiveness. Kurita et al. (Kurita et al., 2020) designed a weight poisoning attack that can preserve the backdoor to the fine-tuned model. They used regularization methods to adjust the model weights, which can reduce the impact of the fine-tuning process on the backdoor. They then modified the embedding vector of triggers to approximate the embedding vector of the target words. Huang et al. (Huang et al., 2023c) believed that backdoor attacks for LLMs required a lot of data and computing resources, reducing the backdoor attacks’ practicality. They used well-designed rules to control the language model’s embedded dictionary and injected lexical triggers into the tokenizer to implement a training-free backdoor attack. Inspired by model editing, Li et al. (Li et al., 2024a) designed a lightweight backdoor editing method for LLMs. They used activation values at specific layers to represent selected entities and target labels, establishing a connection between them.
@@
Data Poisoning Sun et al. (Sun et al., 2023) Static trigger Modify data labels
Wallace et al. (Wallace et al., 2021) Imperceptible trigger Optimize prompts to generate poisoned data
@@
Yang et al. (Yang et al., 2021a) Static trigger Poison the embedding layer
@@ 4.3. Countermeasures of pre-training LLMs
For the various risks during the pre-training phase, we explored countermeasures from two aspects: privacy protection and security defense.
@@ 4.3.1. Privacy protection
As shown in Figure 4, defenders can use two types of countermeasures to mitigate the privacy risks faced in the pre-training scenario: corpora cleaning and privacy pre-training.
Corpora cleaning.
@@
LLMs tend to memorize private information from the training data, resulting in privacy leakage risks in their outputs. Currently, the mainstream method for mitigating such risks involves corpora cleaning (Pan et al., 2020; Hernandez et al., 2022). For example, Subramani et al. (Subramani et al., 2023) and Uzuner et al. (Uzuner et al., 2007) identified texts carrying personal information from datasets and removed them. While Ruch et al. (Ruch et al., 2000) proposed a dictionary-based method that uses predefined rules to identify personal information. Then, some researchers (Achiam et al., 2023) employed neural networks to detect personal information, including OpenAI. Kandpal et al. (Kandpal et al., 2022) and Lee et al. (Lee et al., 2022) noted the significant impact of data duplication on privacy protection. According to experiments, Johnson et al. (Johnson et al., 2020) demonstrated that removing duplicated data and personal information can reduce the risk of LLMs’ privacy leakage. However, cleaning massive corpora presents several challenges. Defenders must keep data utility when removing private information and consider the computational resources of privacy protection methods.
@@
Privacy pre-training. Concerning malicious users (i.e., white-box privacy attackers), model developers can design privacy protection methods from two aspects: the model architecture and the training process. The model architecture determines how knowledge is stored and how the model works during the training and inference phases, which impacts the privacy protection capabilities of LLMs. Jagannatha et al. (Jagannatha et al., 2021) explored privacy leakage in various language models and found that larger models pose higher risks of information leakage. Currently, research on optimizing model architecture for privacy protection is not extensive and can be used as an empirical approach.
@@
Improving the training process can reduce the privacy risk malicious users pose, differential privacy (Habernal, 2021; Yang et al., 2023). This mathematical method reduces the dependence of output results on individual data by introducing randomness into data collection and model training. At first, Abadi et al. (Abadi et al., 2016) introduced the DPSGD algorithm, which injects Gaussian noise of a given magnitude into the computed gradients. Specifically, this method could meet the privacy budget when training models. Li et al. (Li et al., 2021) found that the training process for LLMs is suitable for the DPSGD algorithm. Their experiments demonstrated that non-standard hyperparameters can balance privacy protection and model performance. For the BERT model, Yu et al. (Yu et al., 2023a) proposed a differential privacy-based selective training algorithm, enhancing the performance of the fine-tuned model. Similarly, Igamberdiev et al. (Igamberdiev and Habernal, 2023) adopted a local differential privacy algorithm to pre-train the BART model. To thoroughly eliminate privacy risks, Mattern et al. (Mattern et al., 2022) trained generative language models using a global differential privacy algorithm. They designed a new mismatch loss function and applied natural language instructions to generate high-quality synthetic.

### v2

@@ 3.3. RAG system
The RAG system is a unique method to enhance the performance of LLMs. This technology does not retrain LLMs and is orthogonal to pre-training and fine-tuning processes. As shown in Figure 1, the system constructs external knowledge bases. When given a prompt, the RAG system retrieves its context from the knowledge base and concatenates it, generating a high-quality response. Figure 8 illustrates the details of the RAG system and gives two malicious entities: contributors and users.
@@
    Malicious contributors. Generally, users aim to construct extensive knowledge bases by collecting data from various sources. However, malicious contributors can poison the knowledge base to conduct backdoor and jailbreak attacks. In this case, the adversary can modify the knowledge base but cannot access the inference process.
@@
    Malicious users. The knowledge bases used by the RAG system contain sensitive and valuable information. Therefore, malicious users can design prompts to steal this information, thereby violating the knowledge owners’ privacy. Moreover, malicious users exploit vulnerabilities in the knowledge bases to create jailbreak prompts that can extract the training data. In this case, the adversary can only access the input interfaces of LLMs.
@@ 3.4. Deploy LLMs
Model owners deploy well-trained LLMs to provide specialized services to users. Since LLMs can understand and follow natural language instructions, users can design specific prompts to achieve their goals. This is known as prompt engineering, exemplified by AUTOPROMPT (Shin et al., 2020), Prompt Tuning (Liu et al., 2023f), and P-Tuning (Liu et al., 2021). As shown in Figure 10, model owners provide user access interfaces to minimize privacy and security risks. Therefore, we consider a black-box attacker who aims to induce various risks through prompt design. Subsequently, we categorize these risks by their specificity to LLMs.
@@
    Unique risks for LLMs. In contrast to small-scale models, LLMs have safety guardrails. Thus, malicious users can design prompts to bypass these guardrails and obtain harmful or leaky outputs. Additionally, LLM prompts contain valuable and sensitive information, and malicious users can perform prompt stealing attacks to compromise the privacy of model owners.
@@
    Common risks to all language models. Regarding the knowledge boundaries of language models, malicious users can construct adversarial prompts by adding adversarial perturbations, causing the model to produce meaningless outputs. Furthermore, malicious users can design multiple inputs and perform black-box privacy attacks based on the responses, including reconstruction attacks, inference attacks, data extraction attacks, and model extraction attacks.
@@ 3.5. Deploy LLM-based agents
LLM-based agents combine the robust semantic understanding and reasoning capabilities of LLMs with the advantages of agents in task execution and human-computer interaction, presenting significant potential. These agents can process various tasks and have become a focal point in research. Compared to LLMs, LLM-based agents are autonomous, allowing them to execute tasks independently rather than passively responding to prompts. As shown in Figure 14, many privacy and security threats exist when deploying these agents. For example, a malicious user can query an LLM-based agent with crafted prompts, thereby accessing private data or performing illegal actions. In this context, we consider two malicious entities: users and agents.
@@
    Malicious users. Due to alignment tuning, LLM-based agents possess safety guardrails. Thus, malicious users can perform jailbreak attacks to bypass these agents’ guardrails, thereby obtaining harmful or leaky outputs. Compared to LLMs, LLM-based agents perform more complex operations at query time, such as interacting with other LLM-based agents or open-source websites. These autonomous operations render jailbreak attacks more challenging. In this case, the attacker only has access to the interfaces of LLM-based agents.
@@
    Malicious agents. Before deploying an LLM-based agent, attackers can inject a backdoor into the agent, as discussed in Section 3.2. When processing a trigger request, an infected agent will perform a predefined action to achieve the attacker’s goal. In personal assistant applications, a backdoored agent will covertly send fraudulent text messages to users’ emergency contacts when given a trigger query. Additionally, an LLM-based agent may execute unauthorized interactions with other LLM-based agents, exposing the private information of all users with whom these agents interact. In a network with multiple interacting agents, humans cannot supervise the interactions between agents, resulting in malicious agents contaminating other entities in the network.
@@ 4. The Risks and Countermeasures of Pre-training LLMs
Section 3.1 presents three threat models for pre-training LLMs, and Figure 2 illustrates the details of each adversary. First, we describe the privacy and security risks associated with these threat models. Then, we present existing countermeasures to offer researchers additional defenses.
@@ 4.1. Privacy risks of pre-training LLMs
In this scenario, developers must collect corpora from sources such as books, websites, and code bases. Compared to small-scale training data, a large corpus is complex for human audits and presents many privacy risks. Inevitably, large texts contain PII (e.g., names) and sensitive information (e.g., health records). Kim et al. (Kim et al., 2024) found that the quality of corpora has a significant impact on LLMs’ privacy. Due to their advanced learning abilities, LLMs can output private information when given specific prefixes (Carlini et al., 2022). For example, if one of the training records is ‘Bob’s email is bob08@gmail.com’, then ‘Bob’s email is’ as a prompt will cause the trained LLM to output ‘bob08@gmail.com’. Once private information is leaked, it poses a serious threat to users’ privacy.
The privacy risk posed by malicious users is common to all models. This indicates that pre-trained LLMs may leak private data. Specifically, adversaries can access the model’s parameters and interface to steal training data through white-box privacy attacks. Nasr et al. (Nasr et al., 2023) applied existing data extraction attacks to measure the privacy protection capabilities of open-source models. They found that larger models generally leaked more data. Assuming a white-box adversary, Rahil et al. (Parikh et al., 2022) framed data extraction as an optimization problem and stole four digits in the training data. Subsequently, a more powerful extraction attack was proposed to steal targeted training data (Zhang et al., 2023b). This attack induced model memory by soft prompt tuning and then used loss smoothing to increase the probability of generating the correct suffixes. Through calibrated confidence estimation, the attacker could accurately select the most likely suffix to extract the targeted data.
@@ 4.2. Security risks of pre-training LLMs
Similar to the risk posed by private data, toxic data in corpora also leads to unique and inherent risks in LLMs, as shown in Figure 2. Huang et al. (Huang et al., 2023b) defined toxic data as disrespectful language, including illegal and offensive texts. As Ousidhoum et al. (Ousidhoum et al., 2021) point out, LLMs’ toxicity is due to the toxic content in the training data. Shaikh et al. (Shaikh et al., 2023) explored the Chain-of-Thought (CoT) scenario in LLMs. They found that zero-shot CoT reasoning in sensitive domains increases the model’s likelihood of producing toxic outputs. Deshpande et al. (Deshpande et al., 2023) found that role-playing can increase the probability that LLMs generate toxic content. Additionally, over-detection can degrade the performance of LLMs. Researchers need to address this risk from multiple aspects to minimize the impact of toxic data.
After pre-training, model developers often upload the trained LLMs to the open community for profit. In this case, a security risk is common to all models. Malicious developers can access the training data and manipulate the model’s pre-training process. Specifically, they can compromise models through poison or backdoor attacks, affecting downstream tasks. Once a user deploys a compromised model, the attacker can construct trigger prompts to obtain desired outputs. Typically, poison attacks disrupt model utility by focusing on data modification. Shan et al. (Shan et al., 2023) designed an efficient poison attack against text-to-image models. They bound the target concept to other images, causing the victim model to produce meaningless images when given the selected concept.
@@
Backdoor attacks aim to compromise model integrity by injecting hidden backdoors (Wang et al., 2023a). As demonstrated in Figure 3, the adversary sets a trigger mode and target content. Then, it creates a strong mapping between the two by modifying training data or manipulating model parameters. When given a trigger prompt, the compromised model will produce a predefined backdoor behavior. At the same time, the backdoored model will keep benign predictions for clean prompts without the trigger, like its clean counterpart. Some researchers initially used static texts as triggers in the NLP domain, such as low-frequency words or sentences (Sheng et al., 2022; Yang et al., 2024c). Wallace et al. (Wallace et al., 2021) used a gradient-based optimization algorithm to construct backdoored samples without explicit triggers, improving the concealment of backdoor attacks. Li et al. (Li et al., 2023g) used state-of-the-art LLMs (e.g., ChatGPT) to construct backdoored samples by rewriting text styles, extending the backdoor attack to the semantic level. Zhao et al. (Zhao et al., 2023b) found that mislabeled texts are difficult to bypass human audit and proposed a clean-label backdoor attack for LLMs. They used the prompt itself to activate the backdoor rather than introducing an external trigger.
@@
In addition to designing various trigger modes, researchers manipulate the training process, as demonstrated by Yan et al. (Yan et al., 2023). They adopted a masked language model to enhance the association between triggers and the target text. Most users can download the open BERT model and adapt it to downstream tasks using local data. Du et al. (Du et al., 2023) designed a supervised contrast learning to extract more robust trigger features using various pre-trained models. Kurita et al. (Kurita et al., 2020) designed a weight poisoning attack that can preserve the backdoor in the fine-tuned model. They used regularization methods to adjust the model weights, reducing the impact of the fine-tuning process on the backdoor. They then modified the embedding vector of triggers to approximate the embedding vector of the target words. Huang et al. (Huang et al., 2023c) argued that backdoor attacks for LLMs require extensive data and computing resources, reducing their practicality. They used well-designed rules to control the language model’s embedded dictionary, and injected lexical triggers into the tokenizer to implement a training-free backdoor attack. Inspired by model editing, Li et al. (Li et al., 2024a) designed a lightweight backdoor editing method for LLMs. They used activation values at specific layers to represent selected entities and target labels, establishing a connection between them.
@@
Data Poisoning Wallace et al. (Wallace et al., 2021) Imperceptible trigger Optimize prompts to generate poisoned data
@@ 4.3. Countermeasures of pre-training LLMs
For the various risks in the pre-training scenario, we explored countermeasures from two perspectives: privacy protection and security defense.
@@ 4.3.1. Privacy protection
As shown in Figure 4, defenders can employ two types of countermeasures to mitigate privacy risks in the pre-training scenario: corpora cleaning and privacy pre-training.
@@
Corpora cleaning. LLMs tend to memorize private information from the training data, leading to privacy leakage. Currently, the mainstream method for mitigating such risk involves corpora cleaning (Hernandez et al., 2022). For example, Subramani et al. (Subramani et al., 2023) identified texts carrying PII from datasets and removed them. Ruch et al. (Ruch et al., 2000) proposed a dictionary-based method using predefined rules to identify PII. Meanwhile, Achiam et al. (Achiam et al., 2023) trained meta neural networks to detect PII. Additionally, Kandpal et al. (Kandpal et al., 2022) noted the significant impact of data duplication on privacy protection. Their experiments demonstrated that removing duplicated data and personal information can reduce the risk of LLM privacy leakage. However, cleaning massive corpora presents several challenges. Defenders need to maintain data utility when removing private information, and consider the computational resources required for privacy protection methods.
@@
Privacy pre-training. Regarding malicious users (i.e., white-box privacy attackers), model developers can design privacy protection methods from two perspectives: the model architecture and the training process. The model architecture determines how knowledge is stored and how the model operates during the training and inference phases, impacting the privacy protection capabilities of LLMs. Jagannatha et al. (Jagannatha et al., 2021) explored privacy leakage in various language models and found that larger models pose higher risks of information leakage. Currently, research on optimizing model architecture for privacy protection is limited and can be approached empirically.
@@
Improving the training process can reduce the privacy risk posed by malicious users through differential privacy (Yang et al., 2023). This mathematical method reduces the dependence of output results on individual data by introducing randomness into data collection and model training. Initially, Abadi et al. (Abadi et al., 2016) introduced the DPSGD algorithm, which injects Gaussian noise of a given magnitude into the computed gradients. Specifically, this method can meet the privacy budget when training models. Li et al. (Li et al., 2021) found that the training process for LLMs is suitable for the DPSGD algorithm. Their experiments demonstrated that non-standard hyperparameters can balance privacy protection and model performance. For the BERT model, Yu et al. (Yu et al., 2023a) proposed a differential privacy-based selective training algorithm, enhancing the performance of the fine-tuned model. To thoroughly eliminate privacy risks, Mattern et al. (Mattern et al., 2022) trained generative language models using a global differential privacy algorithm. They designed a new mismatch loss function and applied natural language instructions to collect high-quality synthetic texts.